OpenStudy (anonymous):

use the substitution method to solve the given system of equations. 4x+2y=7 y=5x A.(2/5,2) B.(1/2,5/2) C.(2,10) D.(7/11,35/11)

OpenStudy (smarty):

4x + 2(5x)=7 4x+10x=7 14x=7 x=1/2 Plug that back in, Your answer is B